OpenStudy (anonymous):

What is the derivative of sin x /(1+sin x) ?

OpenStudy (inkyvoyd):

d/dx uv=(u'v-uv')/v^2 u=sin x v=1+sin x

OpenStudy (inkyvoyd):

u'=cos x v'=cos x

OpenStudy (anonymous):

Where I'm confused is what am i supposed to do with [(1+sin x)(cos x) - (sin x)(cos x)] in the numerator?

OpenStudy (inkyvoyd):

expand it, and simplify the entire solution.

OpenStudy (anonymous):

That was my first question

OpenStudy (inkyvoyd):

[(1+sin x)(cos x) - (sin x)(cos x)]=(cos x+sin x cos x-sin x cos x)=(cos x)

OpenStudy (anonymous):

Sorry I saw it just now, duh. Thanks

OpenStudy (inkyvoyd):

np. make sure you put in the denom!

OpenStudy (bradely):

step by step answers posted here http://www.mathskey.com/question2answer/2061/what-is-the-derivative-of-sin-x-1-sin-x ask your calculus homework questions at http://www.mathskey.com/question2answer/calculus and get free math help. all the best